OpenStudy (anonymous):

@halorazer Solve by the linear combination method (with or without multiplication). x + y = 40 0.08x + 0.03y = 1.7 A. (40, 60) B. (20, 20) C. (10, 30) D. (20, –50)

OpenStudy (anonymous):

-12.5(.08x+.03y=1.7) -x-.375y=-21.25 Now we can combine this with x+y=40 x+y=40 -x-.375y=-21.25 0x+.625y=18.75 .625y=18.75 /.625 y=30 Now just put that back into .08x+.03y=1.7 .08x+.9=1.7 -.9 .08x=.8 /.08 x=10 (10,30) So, C.

OpenStudy (anonymous):

THANK YOU SO MUCH! I've been trying to figure out how to do that for like an hour o.0 Thank you! :D

OpenStudy (anonymous):

You're very welcome. :) If you need anymore help, you know where to find me.

OpenStudy (anonymous):

Yep :D